given f(x)={(3,5),(2,4),1,7)} and h(x)={(3,2),(4,3),(1,6)}
find (f+h)(1)

Answer by stanbon(75887)

given f(x)={(3,5),(2,4),(1,7)} and h(x)={(3,2),(4,3),(1,6)}
find (f+h)(1)
---
(f+h)(1) = f(1) + h(1) = 7 + 6 = 13
====================================
Cheers,
Stan H.
=======